### v1

# An investigation of air pollution-induced temperature sensitivity and susceptibility to heat-related hospitalization in the Medicare population

Abstract
Background: With rising temperatures and an aging US population, understanding how to prevent heat-related illness among older Americans will be an increasingly critical objective. Despite biological plausibility, no study to date has investigated how exposure to fine particulate matter air pollution (PM2.5) may contribute to risk of heat-related hospitalization.
Methods: We identified Medicare fee-for-service beneficiaries who experienced a heat-related hospitalization between 2008 and 2016. Using a case-crossover design and fitting Bayesian conditional logistic regression models, we characterized the association between heat-related hospitalization and temperature and PM2.5 exposures. We estimated the relative excess risk due to interaction (RERI) to quantify the additive-scale interaction of simultaneous exposure to heat and PM2.5.
Results: We observed 112,969 heat-related hospitalizations. Fixing PM2.5 at the case day median, increasing temperature from its case day median to the 95th percentile was associated with an odds ratio of 1.045 (95% CI: 1.026, 1.063). Fixing temperature at the case day median and increasing PM2.5 from its median to the 95th percentile was associated with an odds ratio of 1.014 (95% CI: 0.993, 1.037). We estimated the RERI associated with simultaneous median-to-95th percentile increases in temperature and PM2.5 to be 0.032 (0.007, 0.057).
Conclusion: Using nationwide Medicare claims and a self-matched study design, we found evidence supporting synergism between temperature and PM2.5 exposures on the risk of heat-related hospitalization.
Keywords: Heat-related illness; Temperature; Air pollution; Exposure synergism; Case-crossover; Medicare
1 Introduction
Heat is a leading cause of weather-related illness and death, and its impact is expected to intensify with climate change (Ebi et al.,, 2021). Previous research suggests that each year, thousands of deaths in the United States are directly or indirectly attributable to heat exposure (Weinberger et al.,, 2020). Extreme heat events have been associated with increased hospitalizations for conditions that may be triggered or exacerbated by heat, such as respiratory (Anderson et al.,, 2013), renal (Fletcher et al.,, 2012), and cardiovascular (Sun et al.,, 2018) diseases, as well as conditions that occur as a direct result of heat exposure (Layton et al.,, 2020), which are broadly referred to as heat-related illness. Heat-related illness ranges in severity from mild conditions such as heat cramps and syncope to severe, life-threatening conditions like heat stroke. Many of these outcomes are preventable through proactive measures such as avoiding direct sun, maintaining hydration, or temporarily relocating to public cooling centers—actions often prompted by heat alerts (Hopp et al.,, 2018). However, despite these preventive efforts, heat-related emergency department visits in the United States reached record levels during the summer of 2023 (Vaidyanathan,, 2024).
Older adults are particularly vulnerable to heat due to impaired thermoregulation, social isolation, a higher prevalence of comorbidities, and the use of medications that increase heat sensitivity (Khalaj et al.,, 2010; Flynn et al.,, 2005; Layton et al.,, 2020). As the US population ages and extreme heat events become more frequent and intense (Pachauri et al.,, 2014), it will become increasingly essential to understand and implement strategies to protect older Americans from heat-related illness.
Previous studies have found synergistic effects between acute exposure to heat and particulate matter (PM) air pollution on mortality (Stafoggia et al.,, 2008; Li et al.,, 2017; Kioumourtzoglou et al.,, 2016) and on certain cause-specific hospitalizations (Yitshak-Sade et al.,, 2018). However, despite biological plausibility, no study has yet investigated the role of PM in heat-related hospitalizations. Fine particulate matter (PM2.5) is small enough to penetrate the lungs and bloodstream, affecting several organ systems. Evidence from epidemiological and toxicological studies show that acute exposure to PM2.5 can impair vascular function, trigger oxidative stress and inflammation, and increase blood pressure (Pryor et al.,, 2022; Basith et al.,, 2022). These physiological responses may hinder the body’s ability to thermoregulate in hot conditions, potentially increasing the risk of heat-related illness when PM2.5 levels are high. Understanding the impact of PM2.5 on heat-related illness is critical for improving heat alert systems and guiding preparedness efforts.
To address this gap, we investigate the independent and combined effects of temperature and PM2.5 on heat-related hospitalizations among older Medicare fee-for-service (FFS) beneficiaries. Specifically, our analysis focuses on inpatient hospitalizations for heat-related illness during the warm season (June–September) from 2008 to 2016. We employ a case-crossover design and a Bayesian analytic approach to achieve two main objectives: 1) reduce bias from unmeasured, time-invariant confounders (Basu et al.,, 2005; Maclure and Mittleman,, 2000), and 2) provide interpretable inferences using posterior distributions generated by Markov chain Monte Carlo (MCMC) sampling. We model nonlinear associations between the exposures and outcomes, and we assess synergy between heat and PM2.5 by estimating the relative excess risk due to interaction (RERI), an additive-scale interaction quantity.
2 Methods
2.1 Study population and case definition
We assessed records from a 50% random sample of Medicare FFS beneficiaries aged 65 and older who resided in the contiguous United States and were continuously enrolled in Parts A, B, and D for at least 12 months between 2008 and 2016. These records include data on beneficiaries’ sex, age, race/ethnicity, Medicaid eligibility status, ZIP code of residence, inpatient admission and discharge dates, and date of death, if applicable. We identified inpatient hospitalizations with a primary or secondary diagnosis of heat-related illness, based on International Classification of Diseases (ICD; 9th and 10th revisions) diagnosis codes. Detailed ICD-9 and -10 codes are provided in the supplemental materials (Table S1). We refer to the composite of these indications as heat-related hospitalization, or HRH. Our analysis is limited to individuals who experienced an HRH during the warm season (June–September), and we included only each beneficiary’s first HRH during the study period, excluding any subsequent events.
2.2 Exposure assessment
Daily maximum temperature data were obtained from the Parameter-elevation Relationships on Independent Slopes (PRISM) model (PRISM Climate Group,, 2023; Daly et al.,, 2008), which estimates temperature across the contiguous United States using surface station data and geographic factors such as elevation, coastal proximity, and atmospheric conditions (Daly et al.,, 2008). Daily temperature estimates are available at a 4 kilometer (km) resolution. For each ZIP code, we assigned the temperature estimate from the PRISM grid cell whose centroid was nearest to the ZIP code’s population weighted centroid.
Daily PM2.5 concentrations were derived from a validated ensemble model that combines ground monitoring data, chemical transport models, satellite observations, and land-use variables (Di et al.,, 2019, 2025). These estimates, available at a 1 km 1 km resolution, were aggregated to the ZIP code level by averaging grid cell values whose centroids fell within each ZIP code.
For each individual and date, exposure was defined as: 1) the daily maximum temperature on that date and 2) the three-day average PM2.5 concentration, calculated as the mean of the current and two preceding days. We used an extended exposure window for PM2.5 to account for the delayed health effects that might be incurred by air pollution exposures, which has been observed in previous studies (Wang et al.,, 2018; Fernández-Niño et al.,, 2018).
2.3 Statistical analysis
We used a case-crossover design to evaluate the effects of acute heat and PM2.5 exposures among individuals who experienced an HRH, with each individual serving as their own control. The case day was defined as the admission date of an individual’s first HRH during the study period. Control days were assigned bidirectionally within the same calendar month and matched to the same day of the week as the case day, with up to four matched control days per case selected before or after the hospitalization date. To eliminate potentially undue influence from extreme PM2.5 concentration outliers, we calculated the 95th percentile of three-day PM2.5 exposures across all case and control days and excluded observations exceeding this threshold (Di et al.,, 2017; Nethery et al.,, 2023).
We implemented the Bayesian analogue to conditional logistic regression using the brms package (Bürkner,, 2017) to estimate the associations of temperature and PM2.5 with HRH. The primary model incorporated natural cubic splines for temperature and PM2.5, along with a linear interaction term. As a sensitivity analysis, we also applied a tensor product spline approach to enable a more complex, nonlinear interaction between the exposures. Detailed descriptions of both approaches are provided in the supplementary materials.
We estimated the odds of HRH under various temperature and PM2.5 scenarios, reporting posterior means and 95% Bayesian credible intervals (CIs) derived from the posterior distribution samples of the MCMC. While the measures of association from conditional logistic regression models are typically evaluated on a multiplicative scale, interpreting multiplicative interactions can be challenging to interpret in a public health context (Rothman,, 2012). Instead, we assess the interaction between heat and PM2.5 on an additive scale by estimating the relative excess risk due to interaction (RERI) (Knol et al.,, 2007; VanderWeele and Knol,, 2014). The RERI quantifies the increased risk associated with simultaneous exposure to heat and PM2.5 beyond the sum of their individual effects. Because HRH is a rare outcome, the odds ratio-based RERI estimate should closely approximate the conventional RERI, which is a function of relative risks (VanderWeele and Knol,, 2014).
To estimate the RERI for two continuous exposures, we select the median and 95th percentile values from the distributions of temperature and three-day average PM2.5 concentrations on case days as contrast levels. This approach to assessing synergism between exposures has been applied in similar epidemiological studies (Fayyad et al.,, 2024; Josey et al.,, 2023). Additional methodological details are provided in the supplemental materials. Finally, we also report the interactions between temperature and PM2.5 on a multiplicative scale as recommended by VanderWeele and Knol, (2014) and Vandenbroucke et al., (2007), estimating this measure using posterior means and 95% Bayesian CIs from the linear interaction coefficient.
All analyses were completed in R version 4.4.0 (R Core Team,, 2021); code to reproduce these analyses is available at https://github.com/laurenmock/temp_pm25_heat-hosp.
3 Results
3.1 Main Results
Among 29,345,820 Medicare beneficiaries in our sample, we observed at least one heat-related hospitalization for 112,969 unique individuals, who served as cases in our case-crossover study. The characteristics of the individuals who experienced an HRH are detailed in Table 1. Most of the hospitalized individuals were women (68.0%) and white (82.3%), with 22.9% eligible for Medicaid. The mean age at the time of hospitalization was 81.7 years. On case days, the mean maximum daily temperature was 84.6F with the 5th and 95th percentile exposures at 68.7F and 98.4F, respectively. Prior to trimming, the case day distribution of three-day PM2.5 concentrations was heavily right skewed with exposures as high as 120.4 . Case and control days with three-day PM2.5 exposures above 18.4 , the 95th percentile of the distribution, were trimmed; this resulted in a mean PM2.5 concentration of 9.4 on case days. PM2.5 exposure levels were notably high among Black beneficiaries and low among American Indian/Alaska Native beneficiaries (Table 1). The full distributions of temperature and PM2.5 case day exposures are displayed in Figures S1 and S2 in the supplementary materials.
| Case day exposures | |||
| Mean (SD) | |||
| Daily max. | Three-day | ||
| N (%) | temperature (F) | PM2.5 () | |
| Total | 112,969 | 84.6 (9.0) | 9.4 (3.7) |
| Sex | |||
| Male | 35,930 (31.8%) | 84.9 (9.1) | 9.4 (3.7) |
| Female | 77,039 (68.2%) | 84.5 (9.0) | 9.5 (3.7) |
| Age category | |||
| 65-74 | 19,165 (17%) | 85.1 (9.1) | 9.6 (3.7) |
| 75-84 | 45,868 (40.6%) | 84.8 (9.0) | 9.4 (3.7) |
| 85+ | 47,936 (42.4%) | 84.3 (9.0) | 9.4 (3.7) |
| Race/ethnicity | |||
| White | 92,975 (82.3%) | 84.4 (9.0) | 9.3 (3.7) |
| Black/African-American | 13,123 (11.6%) | 85.9 (8.2) | 10.3 (3.5) |
| Asian/Pacific Islander | 2,006 (1.8%) | 82.7 (9.4) | 9.8 (3.6) |
| Hispanic | 3,178 (2.8%) | 87.8 (9.6) | 9.3 (3.4) |
| American Indian/Alaska Native | 650 (0.6%) | 86.6 (10.3) | 7.6 (3.8) |
| Other | 1,037 (0.9%) | 84.4 (9.5) | 9.6 (3.6) |
| Medicaid eligibility | |||
| Ineligible | 87,059 (77.1%) | 84.5 (9.0) | 9.3 (3.7) |
| Eligible | 25,910 (22.9%) | 85.1 (9.1) | 9.9 (3.6) |
Temperature exposure on case days was highest in the South and Southwest regions of the US (Figure 1, left panel), while PM2.5 exposure concentrations were highest in the South, Midwest, and Southern California (Figure 1, middle panel). The rates of HRH per 1,000 Medicare FFS beneficiaries at risk varied substantially across the United States, with the highest rates in the Gulf Coast and Great Plains regions (Figure 1, right panel).
Figure 2 visualizes the independent nonlinear effects of heat and PM2.5 on heat-related hospitalization; we hold PM2.5 at its case day median to examine the effect of heat, and hold temperature at its case day median to examine the effect of PM2.5. Maximum daily temperature exhibited a strong positive association with odds of HRH (Figure 2, left panel). Specifically, when we fix PM2.5 concentrations at 8.9 , increasing the temperature from 85.3F (median) to 98.4F (95th percentile) yields an odds ratio (95% CI) of 1.05 (1.03, 1.06) (Figure 3 and Table S2 in the supplementary materials). Although a positive association was observed between three-day PM2.5 and HRH, credible intervals were wide and covered one across the range of PM2.5 levels (Figure 2, right panel). When we fix temperature at 85.3F, an increase in PM2.5 concentrations from 8.9 (median) to 16.1 (95th percentile) corresponds to an odds ratio (95% CI) of 1.01 (0.99, 1.04) (Figure 3 and Table S2 in the supplementary materials).
We observed an additive interaction between temperature and PM2.5. A simultaneous increase in both temperature and PM2.5 from their median values to their respective 95th percentile quantities was associated with an odds ratio (95% CI) of 1.09 (1.06, 1.12) (Figure 3 and Table S2 in the supplementary materials). Using these odds ratios to approximate the RERI (95% CI), we found that the approximate risk increase was 0.03 (0.01, 0.06) (Figure 3 and Table S2 in the supplementary materials), indicating synergism between the two exposures resulting in excess heat-related hospitalizations. We visualized synergism between these two exposures by comparing the odds of HRH at median exposure levels to the odds of HRH across a grid of exposure values (Figure 4). Temperature is strongly associated with HRH, but on hot days, the odds of HRH also increases with higher concentrations of PM2.5. On cooler days, however, we see a small protective effect of PM2.5.
Although we were primarily interested in assessing the additive interaction with the RERI, we also estimated the interaction on a multiplicative scale. The estimated odds ratio associated with the linear interaction term was 1.00031 (95% CI: 1.00006, 1.00056). Both the additive and multiplicative interaction estimates were positive with credible intervals that did not cover the null value.
3.2 Sensitivity analysis
To investigate the interaction between temperature and PM2.5 further, we developed an additional model using a tensor product to allow for a more flexible, nonlinear interaction between these exposures. This approach resulted in more linear estimates of the independent effects of temperature and PM2.5 compared to the main analysis (Figure S3 in the supplementary materials). The estimated RERI from this model was positive, but the 95% CI included zero and was smaller than the RERI estimated in the main analysis (Figures S4 and S5 and Table S2 in the supplementary materials). The wider credible intervals from the tensor model reflect increased uncertainty in the estimates when allowing for greater flexibility in the interaction structure.
4 Discussion
4.1 Summary and key findings
To our knowledge, this is the first study to investigate the combined (and synergistic) effects of acute exposure to high temperatures and elevated PM2.5 on heat-related hospitalizations. Comparing median and 95th percentile exposure levels, we found a positive association between temperature and HRH (OR: 1.04; 95% CI: 1.03, 1.06). The association between PM2.5 and HRH was also positive, although the credible interval covered the odds ratio scale null value of one. Finally, we identified evidence of an additive interaction between temperature and PM2.5 (RERI: 0.03; 95% CI: 0.01, 0.06). Nevertheless, since temperature and PM2.5 are often correlated (see supplementary Figure S2), we have limited data to assess the effect of elevated PM2.5 at lower temperatures during the warm season.
Our findings offer the first empirical, epidemiological evidence that PM2.5 may increase risk of heat-related hospitalization. In contrast to prior work, which has largely framed temperature as a modifier of the health effects of air pollution (Li et al.,, 2017; Areal et al.,, 2022; Anenberg et al.,, 2020), our focus on explicit heat-related hospitalizations reveals that PM2.5 may also modify the health effects of heat.
4.2 Impact
Heat is the deadliest form of extreme weather in the United States (National Weather Service, 2024b, ), and rates of heat-related illness continue to climb (Vaidyanathan,, 2024; Osborne et al.,, 2023) even though these illnesses are predominantly preventable (Sorensen and Hess,, 2022; Gauer and Meyers,, 2019). Understanding how to protect vulnerable individuals, including older adults, from extreme heat events is a crucial component of climate change adaptation and an urgent public health issue (Vaidyanathan,, 2024).
The National Weather Service issues localized heat alerts to initiate individual- and community-level preparations on unusually warm days (National Weather Service, 2024a, ). Our findings indicate that additional caution is necessary on hot days when PM2.5 levels are high, and it may be useful to consider PM2.5 when issuing heat alerts. However, the efficacy of heat alerts alone in minimizing illness and death is uncertain (Weinberger et al.,, 2018; Toloo et al.,, 2013). Previous research highlights that high-risk individuals must recognize their own vulnerabilities to heat, possess the knowledge and resources necessary to protect themselves, and take action to stay safe (Toloo et al.,, 2013). Interestingly, previous studies report that many older adults recognize age as a risk factor for heat-related illness but do not recognize high temperatures as a threat to their personal health (Widerynski et al.,, 2017). Clinicians could play an important role in helping these individuals understand their risk (Sorensen and Hess,, 2022).
Our results also emphasize the importance of air conditioning, as open windows may allow more ambient particulate matter into homes (Lin et al.,, 2013) and further increase the risk of heat-related illness. These findings also contribute to the growing body of evidence that PM2.5 is linked to a range of negative health effects, underscoring the importance of reducing ambient PM2.5, particularly with a warming planet.
4.3 Biological plausibility
This study supports mounting epidemiological evidence of synergism between temperature and PM2.5. Stafoggia et al., (2008) identified positive but generally non-statistically significant interactions between temperature and PM10, which includes both PM2.5 and larger particles, on mortality, and a meta-analysis by Li et al., (2017) also reported that extreme temperatures modify the effect of PM10 on non-accidental and cardiovascular mortality. Kioumourtzoglou et al., (2016) observed a stronger association between long-term PM2.5 exposure and mortality in warmer cities. A recent study also identified synergism between short-term exposures to temperature and PM2.5 for both respiratory and cardiac hospital admissions (Yitshak-Sade et al.,, 2018).
Simultaneous exposure to heat and PM2.5 can exacerbate heat-related health outcomes through shared pathophysiological mechanisms, including autonomic dysfunction and heightened inflammatory responses. Both heat and PM2.5 cause autonomic dysfunction by increasing sympathetic drive and suppressing parasympathetic activity (Taylor-Clark,, 2020; Yamamoto et al.,, 2007). This imbalance leads to a 4%–7% reduction in heart rate variability and an increase in heart rate by 5–9 beats per minute (Shah et al.,, 2019; Florea and Cohn,, 2014; Jensen and Brabrand,, 2015). In addition, blood pressure increases due to autonomic vasoconstriction from PM2.5 exposure, with each 10 increment correlating with a 9 mm Hg rise in systolic pressure (Dvonch et al.,, 2009). Furthermore, increased sweating via sympathetic activation in response to heat elevates blood pressure by increasing plasma viscosity by 2% per C (Sakurai et al.,, 2004). Thus, the combined cardiovascular stress disrupts hemodynamic regulation, compromising the body’s thermoregulatory function and exacerbating heat stress.
Heat and PM2.5 can also increase systemic inflammatory responses by forming reactive oxygen species within the cells. The resulting oxidative stress activates redox-sensitive transcription factors, driving the production and release of pro-inflammatory mediators such as tumor necrosis factor-alpha (TNF-), interleukin-6 (IL-6), and interleukin-1 (IL-1) (Sangkham et al.,, 2024; Ebi et al.,, 2021). A 10 increase in PM2.5 exposure corresponds to a fivefold rise in these cytokines, and a 2C elevation in ambient temperature can boost their expression by up to fourfold (Hu et al.,, 2024; Park et al.,, 2024). Elevated cytokines also disrupt endothelial homeostasis and increase vascular permeability (Park et al.,, 2024; Pope III et al.,, 2016; Briet et al.,, 2007). The sustained up-regulation of cytokines affects the hypothalamic thermoregulatory center, while impaired vascular permeability decreases the efficacy of peripheral vasodilatory response and increases the risk of heat-related hospital admissions (Vybíral et al.,, 2005).
4.4 Limitations and strengths
We acknowledge that our study has limitations. As with any case-crossover study, the results may not fully generalize to other populations or demographic groups; that is, the group of individuals who experienced a heat-related hospitalization is comprised of a small percentage of all Medicare FFS beneficiaries, and results drawn from this sample may not generalize to the full population. Furthermore, it is likely that PM2.5 partially mediates the relationship between temperature and various health conditions (Tran et al.,, 2023; Ham et al.,, 2024), including heat-related illnesses, so our estimate of the independent effect of temperature may not fully capture more complex pathways involving PM2.5. Consequently, interpreting temperature and PM2.5 as entirely independent exposures may be overly simple, as these exposures may influence each other’s effects on health outcomes in complex ways. We attempted to capture this with an interaction, but this may not sufficiently capture the true relationship between the exposures.
However, this study also has many strengths. We leverage a large and diverse sample of Medicare FFS beneficiaries, enabling robust estimates of risk across a range of environmental conditions. By applying a case-crossover design, the study effectively controls for time-invariant and slow-varying confounders at the individual level, allowing for more precise inferences of acute exposure effects. Moreover, our use of a flexible statistical model that incorporates both additive and interactive effects of temperature and PM2.5 provides valuable insights into the complex, real-world dynamics between these environmental exposures and their impacts on human health. Together, these strengths support the reliability of our findings and their relevance for guiding targeted public health interventions to mitigate the risks of heat-related hospitalizations in vulnerable populations.
4.5 Future work
As this study is the first to provide epidemiological evidence for synergism between the effects of temperature and PM2.5 on heat-related illness, additional work is needed to understand how these exposures interact. It has been well-established that the dangers of heat vary by geographic region and also time of year (Anderson and Bell,, 2011), since individuals adapt behaviorally and biologically to heat. An analysis that accounts for climate type or region may uncover differences in these effects by location. For example, the threat of PM2.5 may be greater in areas where fewer people have access to air conditioning and are thus more likely to be exposed to ambient air pollution on hot days. PM2.5 is also a diverse category of pollutants that includes a wide array of chemical components from various sources, and these components vary widely by location. Evidence suggests that certain components are more harmful than others (Hao et al.,, 2023; Masselot et al.,, 2022; Feng et al.,, 2024). Lastly, since this study focuses on older adults, there is a need to assess potential synergism among younger vulnerable individuals, such as children or adults who work outdoors. This study is a first step in understanding the complex synergism between the effects of temperature and PM2.5 on heat-related illness, and additional research is essential to understand the risk of simultaneous exposure.
Acknowledgments
This work was funded by the National Institutes of Health grants R01AG060232 and 1K01ES032458 and an Amazon Web Services grant on AI/ML for Identifying Social Determinants of Health.
Data availability and reproducibility
To protect the privacy of Medicare beneficiaries, our data usage agreement with CMS prohibits us from sharing patient data publicly. However, interested parties may purchase these data from CMS through the Research Data Assistance Center (ResDAC). The temperature and PM2.5 daily exposure data we used in our analysis are publicly available at https://prism.oregonstate.edu/ and https://doi.org/10.7910/DVN/58C6HG, respectively. Finally, our R analysis scripts are available at https://github.com/laurenmock/temp_pm25_heat-hosp.
Declaration of interests
Soko Setoguchi received research funding from Pfizer Inc., Pfizer Japan, Bristol Myers Squibb, and Daiichi Sankyo and served as a consultant for Pfizer Japan and Merck & Co., Inc.
Supplementary Materials
Appendix A Methods
A.1 ICD codes
Heat-related hospitalizations were identified when the primary or secondary diagnosis code of an inpatient hospitalization indicated heat-related illness. We used the following ICD-9 and ICD-10 codes to define heat-related illness:
| ICD-9 | ICD-10 | |
| Effects of heat and light | 992 | T67 |
| Exposure to excessive natural heat | E900 (excluding E900.1) | X30 |
| Dehydration | 2765 | E86 |
A.2 Main analysis details
We fit natural cubic splines for the effects of temperature and PM2.5 using the ns() function from the splines R package. Each spline had three degrees of freedom, with knot placement performed in an a priori fashion—the default for the ns() function.
A.3 Sensitivity analysis details
As a sensitivity analysis, we also fit a model using a tensor product smooth to more flexibly capture the interaction between temperature and PM2.5. Tensor product smooths combine separate marginal smooth terms for each covariate, allowing each to have its own basis functions and smoothness penalties. Specifically, we implemented a t2-style tensor product smooth from the mgcv R package (Wood et al.,, 2013), which is compatible with Bayesian modeling in brms. Note that the alternative te-style tensor product smooths, which apply penalties differently, are currently not supported in brms. For further details and guidance on implementing tensor product smooths (t2 and te) using mgcv, see the tutorial by Pedersen et al., (2019).
A.4 OR and RERI estimation details
In a case-crossover study, we use conditional logistic regression to assess how temperature () and PM2.5 () exposures affect the odds of heat-related hospitalization. Let indicate the outcome for observation in beneficiary (e.g., for each beneficiary, for the case day and for the control days). We model the log-odds (logit) of the event for a given day’s exposures as:
Here and are flexible functions (e.g., spline terms) capturing potentially nonlinear effects of temperature and PM2.5, and is an interaction term allowing the combined effect of and to deviate from additivity on the log-odds scale. In our primary model, we assume a linear interaction . In a sensitivity analysis we assume is a tensor product of two spline basis functions. The term is a stratum-specific intercept for each beneficiary, which absorbs baseline risk differences between beneficiaries. Because we condition on each beneficiary, these terms are neither estimated nor reported—in other words, the model does not yield an overall baseline probability of hospitalization. Instead, the regression focuses on estimating the relative effect of exposures within each beneficiary. As a result, the exposure coefficients have a conditional interpretation (i.e., comparing odds of the event between two exposure levels for the same beneficiary) rather than a population-average effect.
The fitted model yields estimates that resemble the log-odds of heat-related hospitalization for any combination of temperature and PM2.5 values for a given beneficiary . Despite the similarities, the fitted values are not actual predictions of the log-odds given the design of the case-crossover study. However, we can still estimate the odds ratio for the outcome from these quantities by contrasting the fitted values of two exposure scenarios. Supposing is the same in both contrasts, exponentiating this difference yields an identifiable odds ratio (OR). For example, consider the distributions of observed case day temperature and PM2.5 exposures. Let be the median temperature and be a higher temperature (e.g., the 95th percentile), and similarly define and for PM2.5. The effect of a higher temperature (vs. median temperature) on hospitalization odds, at a common PM2.5 level, can be expressed as:
Without needing to predict , this quantity is identifiable and estimated by
where and are fit with conditional logistic regression. Similarly, we can examine the odds of a heat-related hospitalization attributable to PM2.5 by estimating
with the plugin estimate
Like and , is fit with conditional logistic regression. Finally, we can also examine the joint effect of high temperature and high PM2.5 by comparing the odds of an event where both exposures are high to a baseline odds when both exposures are at median levels. This quantity is identified as
which is estimated with
To quantify departure from additive effects from predictions on a relative risk scale, we can compute the relative excess risk due to interaction (RERI). In our context, we can define exposures as “high” vs “baseline” (e.g. vs , and vs ) and use the odds ratios as approximations for the relative risk predictions. The RERI is then computed using already estimated values with
Appendix B Additional results
| Main analysis (splines) | Sensitivity analysis (tensor) | |||
| Posterior | Posterior | |||
| Estimate (95% CI) | probability | Estimate (95% CI) | probability | |
| OR for heat (OR10) | 1.045 (1.026, 1.063) | <0.001 | 1.054 (1.027, 1.097) | <0.001 |
| OR for PM2.5 (OR01) | 1.014 (0.993, 1.037) | 0.098 | 1.032 (0.977, 1.101) | 0.116 |
| OR for heat & PM2.5 (OR11) | 1.091 (1.061, 1.124) | <0.001 | 1.106 (1.040, 1.194) | 0.002 |
| RERI | 0.032 (0.007, 0.057) | 0.006 | 0.019 (-0.028, 0.068) | 0.180 |
B.1 Sensitivity analysis results